# DeepGesture: A conversational gesture synthesis system based on emotions and semantics

### v2

@@ Abstract.
In this work, we present DeepGesture, a diffusion-based gesture synthesis framework for generating expressive co-speech gestures conditioned on multimodal signals—text, speech, emotion, and seed motion. Built upon the DiffuseStyleGesture model, DeepGesture introduces novel architectural enhancements that improve semantic alignment and emotional expressiveness in generated gestures. Specifically, we integrate fast text transcriptions as semantic conditioning and implement emotion-guided classifier-free diffusion to support controllable gesture generation across affective states. To visualize results, we implement a full rendering pipeline in Unity based on BVH output from the model. Evaluation on the ZeroEGGS dataset shows that DeepGesture produces gestures with improved human-likeness and contextual appropriateness. Our system supports interpolation between emotional states and demonstrates generalization to out-of-distribution speech, including synthetic voices—marking a step forward toward fully multimodal, emotionally aware digital humans. Our paper is featured on the  [^1] https://deepgesture.github.iohomepage
@@
copyright: acmlicensedjournalyear: 2026doi: 10.1145/XXXXXXX.XXXXXXXconference: ACM SIGGRAPH Conference on Computer Graphics and Interactive Techniques; July 19–23, 2026; Los Angeles, CA, USAisbn: 978-1-4503-XXXX-X/26/07ccs: Computing methodologies Animationccs: Computing methodologies Natural language processingccs: Computing methodologies Neural networks
@@ 1.2. Problem Statement
Figure 3. A gesture sequence: the first $N$ frames are used as seed gesture $\mathbf{s}$, and the remaining $M$ frames are to be predicted
@@ 2.2.3. Deep Learning Methods
  • HA2G (Liu et al., 2022a) builds a hierarchical Transformer-based model to learn multi-level correlations between speech and gestures, from local to global features.
@@
  • Cascaded Motion Network (CaMN) (Liu et al., 2022b) introduces the BEAT dataset and a waterfall-like model. Speaker information, emotion labels, text, speech, and gesture features are processed through layers to extract latent vectors. In the fusion stage, CaMN combines features sequentially: speaker and emotion features are merged first, followed by integration with latent vectors of text, speech, and gestures.
@@
Implicit generative models learn the data distribution without explicitly modeling the probability density function $p(\mathbf{x})$. Instead of directly computing $p(\mathbf{x})$, the model learns a mapping $G_{\theta}:\mathcal{Z}\to\mathcal{X}$ by matching the distributions between real and generated data $\mathbf{x}=G_{\theta}(\mathbf{z}),\quad\mathbf{z}\sim p_{z}(\mathbf{z})$. Here, $\mathcal{Z}$ denotes the input noise space, typically with a simple distribution $p_{z}$ (e.g., Gaussian, Uniform), and $\mathcal{X}$ is the space of real data, which in this case is gesture sequences $\mathbf{x}$.
@@ 3. PROPOSED METHOD
Figure 6. Overview of the DeepGesture model
The proposed DeepGesture model is based on the DiffuseStyleGesture model (Yang et al., 2023a), which applies the Diffusion model (Ho et al., 2020) with conditional guidance (Ho and Salimans, 2022) (Classifier-Free Diffusion Guidance) to control features during the denoising process.
@@
The model’s objective is to learn parameters $\theta$ of the generative function $G_{\theta}$ with inputs being the noisy gesture matrix $\mathbf{x}_{t}\in\mathbb{R}^{1:M\times D}$, timestep $t$, and condition $c$. An overview of the proposed DeepGesture model is illustrated in Figure 6. As with standard diffusion models, it includes two processes: the diffusion process $q$ and the denoising process $p_{\theta}$ with weights $\theta$. The 1. Preprocessing stage will be presented in Subsection 4.2.
@@ 3.1. Feature Processing Stage
[…]
@@
[…]
@@ 3.2. Feature Extraction Stage
Figure 7. The Model Architecture in DeepGesture
@@ 3.7. Training Procedure
Algorithm 1 Training in DeepGesture
@@
Algorithm 1 trains the DeepGesture model by first computing the required values and hyper-parameters – $\gamma$, $\sqrt{\alpha_{t}}$, $\sqrt{1-\alpha_{t}}$, $\sqrt{\bar{\alpha}_{t}}$, and $\bm{\epsilon}_{t}$ – for every timestep $t$ (1 … $T$). The initial label $\mathbf{x}_{0}$, representing the ground-truth gesture, is drawn from the normalized data distribution. Random Bernoulli masks $c_{1}$ and $c_{2}$ emulate different conditions (gesture, emotion, speech, or text), with one mask possibly lacking emotion information. Noise is then added to create the noisy gesture $\mathbf{x}_{t}$. A timestep $t$ is sampled uniformly, and $\mathbf{x}_{t}$ with the masks is fed into the model to predict the original gesture sequence as a weighted combination of conditional outputs. The Huber loss between ground-truth and prediction is used to update $\theta$. This cycle repeats until the model converges, yielding the optimal parameters $\theta^{\prime}$.
@@ 3.8. Sampling Process
Algorithm 2 Sampling in DeepGesture
@@
    For each timestep, take $t$ sequentially from $[T,\dots,1]$.
@@ 4.2. Data Preprocessing
Speech data: $\mathbf{a}_{\text{raw}}\in\mathbb{R}^{\text{length }}$ is raw speech sampled at 16000 Hz, trimmed into segments $\mathbf{a}\in\mathbb{R}^{64000}$ corresponding to 4 seconds. The paper uses ffmpeg-normalize to normalize volume to a level lower than the original.
@@ 4.3. Training Process
The training process is implemented in the open-source repository: Github/DeepGesture  [^2] https://github.com/DeepGesture/DeepGesture.
@@ 5.1.1. Human Perception Evaluation
Within the GENEA consortium (Generation and Evaluation of Non-verbal Behaviour for Embodied Agents), evaluators are recruited via Prolific. Participants watch videos rendered by different models and assign ratings of Left Better, Equal, or Right Better. These responses are converted into numerical scores of $-1$, 00, and $1$, respectively, and updated via the Elo rating system.
@@ 5.1.2. Objective Metric Evaluation
(10) $\text{FGD}=\left\|\hat{\mu}-\mu\right\|^{2}+\operatorname{Tr}\left(\Sigma+\hat{\Sigma}-2\sqrt{\Sigma\hat{\Sigma}}\right)$
@@ 5.2.1. User Study Results
Due to the high cost and resource requirements of human evaluations, this work reuses the subjective evaluation results from DiffuseStyleGesture (Yang et al., 2023a) as a baseline. The proposed DeepGesture model is not included in the subjective study.
@@ 5.2.2. Quantitative Results
DeepGesture (Feature D=1141) 2.058 9465.546
DeepGesture (Rotations) 3.513 9519.129
Table 3. FGD results for DeepGesture on $\mathbf{x}^{1:M\times D}$
@@
  • Rotations: Joint rotations are extracted, resulting in $D=225$ features ($75$ joints $\times$ $3$ rotation angles).
@@ 6. STRENGTHS AND LIMITATIONS
The proposed DeepGesture model introduces several key advantages that advance the development of more natural and expressive human-machine interaction systems. Nevertheless, certain limitations remain and represent important directions for future work.
@@
  • High realism: DeepGesture generates gestures that exhibit strong human-likeness, closely mirroring the timing and rhythm of natural speech. The model effectively synchronizes gestures with both the emotional tone and semantic content of speech.
  • Robust generalization: Leveraging the denoising capabilities of diffusion models, DeepGesture can generate gestures for speech and emotional contexts not encountered during training, indicating strong potential for real-world deployment across varied domains.
@@ 7.1. Achieved Results
This work presents DeepGesture, a gesture generation model that achieves high realism and naturalness through diffusion-based modeling. A primary contribution is its precise synchronization between generated gestures and the emotional content of speech, extending generalization beyond the training distribution. The model demonstrates the ability to produce context-aware gestures even in low-probability speech scenarios.
@@ 7.2. Future Work
Several promising directions can further improve DeepGesture and expand its applicability in real-world settings:
  • Real-time inference optimization: Future efforts will focus on transforming DeepGesture into a real-time gesture generation system by reducing dependencies on offline rendering tools (e.g., Unity) and optimizing latency for interactive applications.
@@ 7.3. Closing Remarks
Through experimental validation and qualitative analysis, the DeepGesture model – extending DiffuseStyleGesture-demonstrates the ability to generate realistic gestures for both in – distribution and out-of-distribution speech, including synthetic voices such as that of Steve Jobs (see Appendix B). This highlights the promise of diffusion-based approaches in modeling complex, expressive, and low-frequency gesture behaviors.
@@ A.3. Conversion from Euler Angles to Quaternions
  • $c_{\alpha}=\cos\left(\frac{\alpha}{2}\right),\quad s_{\alpha}=\sin\left(\frac{\alpha}{2}\right)$
  • $c_{\beta}=\cos\left(\frac{\beta}{2}\right),\quad s_{\beta}=\sin\left(\frac{\beta}{2}\right)$
  • $c_{\gamma}=\cos\left(\frac{\gamma}{2}\right),\quad s_{\gamma}=\sin\left(\frac{\gamma}{2}\right)$

### v3

# A conversational gesture synthesis system based on emotions and semantics

@@ Abstract.
In this work, we present OHGesture, a diffusion-based gesture synthesis framework for generating expressive co-speech gestures conditioned on multimodal signals - text, speech, emotion, and seed motion. Built upon the DiffuseStyleGesture model, OHGesture introduces novel architectural enhancements that improve semantic alignment and emotional expressiveness in generated gestures. Specifically, we integrate fast text transcriptions as semantic conditioning and implement emotion-guided classifier-free diffusion to support controllable gesture generation across affective states. To visualize results, we implement a full rendering pipeline in Unity based on BVH output from the model. Evaluation on the ZeroEGGS dataset shows that OHGesture produces gestures with improved human-likeness and contextual appropriateness. Our system supports interpolation between emotional states and demonstrates generalization to out-of-distribution speech, including synthetic voices - marking a step forward toward fully multimodal, emotionally aware digital humans.  [^1] This manuscript is provided for preview purposes only and has not been published.
@@
$c$s: Computing methodologies Animationccs: Computing methodologies Natural language processingccs: Computing methodologies Neural networks
@@ 1.2. Problem Statement
Figure 3. A gesture sequence: the first NN frames are used as seed gesture $\mathbf{s}$, and the remaining $M$ frames are to be predicted
@@ 2.2.3. Deep Learning Methods
  • HA2G (Liu et al., 2022b) builds a hierarchical Transformer-based model to learn multi-level correlations between speech and gestures, from local to global features.
@@
  • Cascaded Motion Network (CaMN) (Liu et al., 2022a) introduces the BEAT dataset and a waterfall-like model. Speaker information, emotion labels, text, speech, and gesture features are processed through layers to extract latent vectors. In the fusion stage, CaMN combines features sequentially: speaker and emotion features are merged first, followed by integration with latent vectors of text, speech, and gestures.
@@
Implicit generative models learn the data distribution without explicitly modeling the probability density function $p(\mathbf{x})$. Instead of directly computing $p(\mathbf{x})$, the model learns a mapping $G_{\theta}:\mathcal{Z}\to\mathcal{X}$ by matching the distributions between real and generated data $\mathbf{x}=G_{\theta}(\mathbf{z}),\qquad\mathbf{z}\sim p_{z}(\mathbf{z})$. Here, $\mathcal{Z}$ denotes the input noise space, typically with a simple distribution $p_{z}$ (e.g., Gaussian, Uniform), and $\mathcal{X}$ is the space of real data, which in this case is gesture sequences $\mathbf{x}$.
@@ 3. PROPOSED METHOD
Figure 6. Overview of the OHGesture model
The proposed OHGesture model is based on the DiffuseStyleGesture model (Yang et al., 2023a), which applies the Diffusion model (Ho et al., 2020) with conditional guidance (Ho and Salimans, 2022) (Classifier-Free Diffusion Guidance) to control features during the denoising process.
@@
The model’s objective is to learn parameters $\theta$ of the generative function $G_{\theta}$ with inputs being the noisy gesture matrix $\mathbf{x}_{t}\in\mathbb{R}^{1:M\times D}$, timestep $t$, and condition $c$. An overview of the proposed OHGesture model is illustrated in Figure 6. As with standard diffusion models, it includes two processes: the diffusion process $q$ and the denoising process $p_{\theta}$ with weights $\theta$. The 1. Preprocessing stage will be presented in Subsection 4.2.
@@ 3.1. Feature Processing Stage
  • Speech $\mathbf{a}\in\mathbb{R}^{1:M\times 1024}$: All speech data in ‘.wav‘ format is downsampled to $16~\mathrm{kHz}$. The speech corresponding to the gesture segment (4 seconds) is extracted as a waveform vector $\mathbf{a}\in\mathbb{R}^{64000}$. Following DiffuseStyleGesture, this work uses the pre-trained WavLM Large model (Chen et al., 2022) to embed the raw waveform into a high-dimensional vector representing acoustic features. Linear interpolation is then applied to temporally align the latent features from WavLM at $20~\text{fps}$, resulting in the speech matrix $\mathbf{a}\in\mathbb{R}^{1:M\times 1024}$.
@@
  • Ground Truth Gesture $\mathbf{x}_{0}\in\mathbb{R}^{1:M\times D}$: This is the ground-truth gesture sequence of $M=80$ frames (corresponding to 4 seconds at $20~\text{fps}$), which is preprocessed to form the matrix $\mathbf{x}_{0}\in\mathbb{R}^{1:M\times D}$.
@@ 3.2. Feature Extraction Stage
Figure 7. The Model Architecture in OHGesture
@@ 3.7. Training Procedure
Algorithm 1 Training in OHGesture
@@
Algorithm 1 trains the OHGesture model by first computing the required values and hyper-parameters – $\gamma$, $\sqrt{\alpha_{t}}$, $\sqrt{1-\alpha_{t}}$, $\sqrt{\bar{\alpha}_{t}}$, and $\bm{\epsilon}_{t}$ – for every timestep $t$ (1 … $T$). The initial label $\mathbf{x}_{0}$, representing the ground-truth gesture, is drawn from the normalized data distribution. Random Bernoulli masks $c_{1}$ and $c_{2}$ emulate different conditions (gesture, emotion, speech, or text), with one mask possibly lacking emotion information. Noise is then added to create the noisy gesture $\mathbf{x}_{t}$. A timestep $t$ is sampled uniformly, and $\mathbf{x}_{t}$ with the masks is fed into the model to predict the original gesture sequence as a weighted combination of conditional outputs. The Huber loss between ground-truth and prediction is used to update $\theta$. This cycle repeats until the model converges, yielding the optimal parameters $\theta^{\prime}$.
@@ 3.8. Sampling Process
Algorithm 2 Sampling in OHGesture
@@
    For each timestep, take $t$ sequentially from $T,\dots,1]$.
@@ 4.2. Data Preprocessing
[…]
@@ 4.3. Training Process
The training process is implemented in the open-source repository: Github/OHGesture  [^2] https://github.com/OHGesture/OHGesture.
@@ 5.1.1. Human Perception Evaluation
Within the GENEA consortium (Generation and Evaluation of Non-verbal Behaviour for Embodied Agents), evaluators are recruited via Prolific. Participants watch videos rendered by different models and assign ratings of Left Better, Equal, or Right Better. These responses are converted into numerical scores of 1-1, 0, and $1$, respectively, and updated via the Elo rating system.
@@ 5.1.2. Objective Metric Evaluation
[…]
@@ 5.2.1. User Study Results
Due to the high cost and resource requirements of human evaluations, this work reuses the subjective evaluation results from DiffuseStyleGesture (Yang et al., 2023a) as a baseline. The proposed OHGesture model is not included in the subjective study.
@@ 5.2.2. Quantitative Results
OHGesture (Feature D=1141) 2.058 9465.546
OHGesture (Rotations) 3.513 9519.129
Table 3. FGD results for OHGesture on $\mathbf{x}^{1:M\times D}$
@@
  • Rotations: Joint rotations are extracted, resulting in $D=225$ features ($75$ joints $\times$ 33 rotation angles).
@@ 6. STRENGTHS AND LIMITATIONS
The proposed OHGesture model introduces several key advantages that advance the development of more natural and expressive human-machine interaction systems. Nevertheless, certain limitations remain and represent important directions for future work.
@@
  • High realism: OHGesture generates gestures that exhibit strong human-likeness, closely mirroring the timing and rhythm of natural speech. The model effectively synchronizes gestures with both the emotional tone and semantic content of speech.
  • Robust generalization: Leveraging the denoising capabilities of diffusion models, OHGesture can generate gestures for speech and emotional contexts not encountered during training, indicating strong potential for real-world deployment across varied domains.
@@ 7.1. Achieved Results
This work presents OHGesture, a gesture generation model that achieves high realism and naturalness through diffusion-based modeling. A primary contribution is its precise synchronization between generated gestures and the emotional content of speech, extending generalization beyond the training distribution. The model demonstrates the ability to produce context-aware gestures even in low-probability speech scenarios.
@@ 7.2. Future Work
Several promising directions can further improve OHGesture and expand its applicability in real-world settings:
  • Real-time inference optimization: Future efforts will focus on transforming OHGesture into a real-time gesture generation system by reducing dependencies on offline rendering tools (e.g., Unity) and optimizing latency for interactive applications.
@@ 7.3. Closing Remarks
Through experimental validation and qualitative analysis, the OHGesture model – extending DiffuseStyleGesture-demonstrates the ability to generate realistic gestures for both in – distribution and out-of-distribution speech, including synthetic voices such as that of Steve Jobs (see Appendix B). This highlights the promise of diffusion-based approaches in modeling complex, expressive, and low-frequency gesture behaviors.
@@ A.3. Conversion from Euler Angles to Quaternions
  • $c_{\alpha}=\cos\left(\frac{\alpha}{2}\right),\qquad s_{\alpha}=\sin\left(\frac{\alpha}{2}\right)$
  • $c_{\beta}=\cos\left(\frac{\beta}{2}\right),\qquad s_{\beta}=\sin\left(\frac{\beta}{2}\right)$
  • $c_{\gamma}=\cos\left(\frac{\gamma}{2}\right),\qquad s_{\gamma}=\sin\left(\frac{\gamma}{2}\right)$